Simplify: cos(4x) cos(2x) - sin(2x) sin(4x)
Found 2 solutions by ikleyn, solve_for_x:
Answer by ikleyn(53765)

.
= cos(6x).

Apply the formula cos(a)*cos(b) - sin(a)*sin(b) = cos(a+b).


Answer by solve_for_x(190)

Using the formula:

cos(a)cos(b) - sin(a)(sin(b) = cos(a + b)

the expression can be simplified as

cos(4x)cos(2x) - sin(2x)sin(4x) = cos(2x + 4x) = cos(6x)